OpenStudy (anonymous):

2 over 3 k + 6 = −12

OpenStudy (anonymous):

2/3k + 6=-12 -6 - 6 2/3k = -18 ---- ----- 2/3 2/3 you can do the rest;) k = ?

OpenStudy (anonymous):

i dont know lol help

OpenStudy (anonymous):

-18 divided by 2/3? Okay one sec...

OpenStudy (anonymous):

k= -27! Hope it helped you got the work;)

OpenStudy (anonymous):

thanks ;)

OpenStudy (anonymous):

No problem! Hope it helped! :)